OpenStudy (anonymous):

The displacement (in centimeters) of a particle moving back and forth along a straight line is given by the equation of motion s = 3 sin πt + 2 cos πt, where t is measured in seconds. (Round your answers to two decimal places.) (a) Find the average velocity during each time period. (i) [1, 2]

OpenStudy (anonymous):

can someone plz help me

OpenStudy (stamp):

\[s'(t)=v(t)\]

OpenStudy (stamp):

\[v=\frac{\Delta s}{\Delta t}\]

OpenStudy (stamp):

\[s(t)=3sin(\pi t)+2cos(\pi t)\]\[s'(t)=v(t)\]\[s'(t)=3\pi cos(\pi t)-2\pi sin(\pi t)\]\[v(t)=3\pi cos(\pi t)-2\pi sin(\pi t)\] \[v(1)=3\pi cos(\pi*1)-2\pi sin(\pi*1)\]\[v(2)=3\pi cos(\pi*2)-2\pi sin(\pi*2)\]